MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2006/08/03
OBSERVER: K.Ohnishi
CONDITION:
COMMENT: Alart on the second caustic crossing at OGRE-2006-BLG-277 ( gb10-7 (x,y)=(817.339,2471.795) ). At the DARK, 60sec dark, D1245-60 at the start and D1251-60 at the end is OK.. At the 300sec dark, D1246 and D1247 is OK. D1244 exits the transvers line in chip 6-10. In D1249, chip7 is very bad. One possiblity at the bad DARK is the saturation due to dawn just before fram.
Compressor gas pressure: MPa
Expected Observation Time: START 18:47 END 07:00


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB3128-3149221478820
LMC561-5622216300
SMC156-156192700
ALERT-000
GRB-000
TOTAL---2517717820 (04h57m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
555522544222244
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
44444444440


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
111111222212
Field13Field14Field15Field16
2200


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
111111111000


SPECIAL TARGET NAME

CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
18:30waitingcloudy and dawn
19:30waitingcloudly
20:30waitingcloudly
21:30waitingcloudly
22:30waitingcloudly
23:30observingfine
00:30observingfine
01:30observingfine
02:30observingfine
03:30observingfine
04:30observingfine
05:30observingfine
06:30observingfine
06:50waiting@stop due to dawn


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day